# Computational Modeling for Personalized Transcranial Electrical Stimulation: Theory, Tools, and Applications

### v1

@@ 2.3 Simulating electrical field distribution induced by tES
where $x$, $y$ and $z$ represent three orthogonal directions, and $w_{x}$, $w_{y}$, and $w_{z}$ denote the weights assigned to the respective orthogonal directions. And because tACS usually has low frequency, a quasi-static approximation can be employed to simulate it. In most cases, we can derive the electric field at peak currents from computational models. The temporal variations of the current merely scale the field without altering its distribution within the brain. When multiple channels are out of phase, simulations should be based on temporal dynamics. Saturnino et al. modeled several in-phase and anti-phase tACS scenarios, finding that they produced complex differences in temporospatial stimulation patterns [67].
tTIS employs a distinct mechanism to stimulate the brain. Two-pair tTIS utilizes two pairs of electrodes, while HD-tTIS employs an array of electrodes for stimulation [68]. Both types of stimulation are designed to stimulate the brain by the low frequency envelope formed by high frequency current. The effect of tTIS can be evaluated by the modulation depth, which is defined as the difference between the maximal and minimal value of the envelope. As models disregard the possible effects of high-frequency stimulation on neurons, they calculate the distinct frequency electric field in the same manner as tACS (Eq. (1), (2)). The envelope field with the desired direction can be calculated by the formula proposed by Grossman et al. [20], shown in Eq. (4).
@@
where $\vec{E}_{1}(\vec{r})$ and $\vec{E}_{2}(\vec{r})$ are the first and second distinct fields at the location $\vec{r}$ and $\vec{n}$ is an unit vector along the desired direction. And Huang et al. provided a comprehensive analysis of the equations governing various forms of tTIS [69, 68]. They showed the formulae for the modulation depth along the maximum direction. It is straightforward to find that Eq. (4) is maximized if and only if $\vec{n}$ falls onto the plane spanned by $\vec{E}_{1}$ and $\vec{E}_{2}$. Under such premise, the maximal modulation depth equation can be formulated as follows.
@@ 2.4 Software for constructing personalized head models
In addition to general FEM platforms, several software packages have been specifically developed for brain stimulation modeling, providing end-to-end solutions for the construction and simulation of personalized head models (see Table LABEL:tab:forward&optim). These include ROAST [70], SimNIBS [66, 71], SCIRun [72], COMETS [73], and SimBio [74]. Among these, ROAST and SimNIBS are the most widely adopted open-source tools in the field. In the following sections, we introduce the computational frameworks underpinning ROAST and SimNIBS, and provide a comparative analysis highlighting their respective strengths, limitations, and suitable application scenarios.
SimNIBS 2.1: SimNIBS 2.1 provides two distinct pipelines for head segmentation and mesh generation: headreco and mri2mesh. Among these, headreco is recognized for its significantly faster processing speed, typically completing within 2 hours, while mri2mesh requires approximately 10 hours to achieve comparable results, according to the official documentation. The mri2mesh pipeline relies on FreeSurfer [75] (version 5.3.0 or newer) and FSL [76] (version 5.0.5 or newer), and thus is not compatible with the Windows operating system. In contrast, headreco utilizes voxel segmentations generated by SPM12 combined with morphological operations, and can run seamlessly across Windows, Linux, and macOS platforms. Notably, headreco supports the use of CAT12 to reconstruct the mid-thickness gray matter (GM) surface, enabling more anatomically accurate individualized head models. Both headreco and mri2mesh share a similar approach to converting voxel segmentations into volume meshes, as described in Windhoff et al. [77].
In terms of segmentation accuracy, MRI-based evaluations reveal that headreco and mri2mesh achieve comparable performance in Dice similarity scores. However, headreco demonstrates a lower Hausdorff distance, indicating more precise delineation of the skull boundaries. Additionally, incorporating a T2-weighted (T2w) image into the segmentation pipeline generally improves the robustness and quality of results [78, 79, 80]. Specifically, the inclusion of T2w images reduces inter-subject variability and mitigates instances of gross segmentation errors, due to the enhanced contrast between skull and cerebrospinal fluid (CSF) compared to T1-weighted (T1w) images.
Regarding mesh quality, both headreco (with CAT12 integration) and mri2mesh produce head meshes of comparable geometric fidelity. The quality of finite element meshes is typically assessed using metrics such as the Aspect Ratio ($\rho$), Volume-to-Edge Ratio ($\eta$), and Volume-to-Surface Ratio ($\tau$), where lower values indicate more regular and well-shaped tetrahedra. Comparative analyses reveal no substantial differences in mesh quality between the two pipelines, likely attributable to their similar surface generation and volume meshing strategies. Interestingly, finer anatomical modeling of cortical folding, especially detailed sulcal structures, slightly increases the occurrence of lower-quality tetrahedra. This effect is observed across segmentation methods, whether employing FSL BET2 and FreeSurfer for mri2mesh, or SPM12 and CAT12 for headreco.
SimNIBS 4: In the latest versions of SimNIBS, the CHARM pipeline has replaced earlier tools such as headreco and mri2mesh, offering improved head segmentation performance. CHARM automatically segments fifteen different head tissues from magnetic resonance (MR) images. A key advantage of CHARM is its robustness to variability in input scans, allowing it to be applied directly to clinical or research data acquired using different scanners, sequences, or imaging protocols. Comparative evaluations using Dice coefficient and modified Hausdorff distance metrics demonstrate that CHARM outperforms headreco and ROAST in simplified segmentation tasks involving the five primary head tissue classes. Moreover, CHARM achieves consistently high segmentation accuracy across all fifteen tissue types, with higher performance observed for larger anatomical structures and slightly lower accuracy for smaller structures. Importantly, CHARM maintains robust segmentation quality even on lower-quality clinical datasets, highlighting its practical applicability across a wide range of imaging conditions.
ROAST: ROAST performs head segmentation by leveraging SPM12, similar to headreco, supplemented with additional post-processing operations. After obtaining the segmentation, ROAST employs Iso2mesh [81], a free MATLAB/Octave-based toolbox for mesh generation and processing, to create tetrahedral meshes. Iso2mesh is capable of generating meshes from surfaces as well as from 3D binary or grayscale volumetric images such as segmented MRI or CT scans. When comparing automated segmentation results to manual annotations, ROAST demonstrates superior performance in segmenting WM, GM, and CSF compared to headreco with CAT12. It shows comparable performance to headreco for skull and air segmentations but exhibits limitations in accurately delineating the scalp region. Additionally, mri2mesh performs worse than headreco in terms of absolute volumetric segmentation differences across multiple tissue types, further underscoring the relative advantages of ROAST and headreco.
Table 2: Software for forward modeling and inverse optimization. ( Only display optimization methods with official implementation, see Section 3.2 for more detailed information )
Method Forward Modeling / Inverse Optimization Website
@@
SimNIBS [82] Both https://simnibs.github.io/simnibs/build/html/index.html
ROAST [70] Both https://github.com/andypotatohy/roast
FreeSurfer [75] Modeling https://surfer.nmr.mgh.harvard.edu/
Brainstorm [83] Modeling https://neuroimage.usc.edu/brainstorm/
@@
DUNEuro [84] Modeling https://www.medizin.uni-muenster.de/duneuro/forschung.html
Iso2mesh [81] Modeling https://iso2mesh.sourceforge.net/cgi-bin/index.cgi
SCIRun [72] Modeling https://sci.utah.edu/software/scirun.html
COMETS [73] Modeling http://cone.hanyang.ac.kr/$b$s/page.php?pa_id=detail6
SimBio [73] Modeling https://simbio.com/
Neurophet tES LAB [85] Modeling https://brainbox-neuro.com/products/neurophet-tes-lab
COMSOL [86] Modeling https://www.comsol.com
Genetic Algorithm [87] Optimization hhttps://zenodo.org/records/5907212
@@
MOVEA [22] Optimization https://github.com/ncclab-sustech/MOVEA
2.5 Application for forward modeling
Forward modeling has proven to be an effective approach for simulating current conduction within the brain and is widely employed in various research domains [12, 88, 89, 90]. The primary applications of forward modeling in research involve targeted electric field simulations and the investigation of neuromodulatory mechanisms (see Table LABEL:tab:application1). Targeted simulations are particularly useful as validation tools before conducting experiments. Many studies utilize open-source head models to compute electric field distributions prior to experimentation, with the goal of assessing whether the stimulation paradigm effectively targets the intended brain regions.
For instance, Luo et al. assessed influence of tACS on the hippocampus and amygdala using real sEEG signals and used SimNIBS to simulate the effects of it. They found that a low-intensity electric field could modulate the activity of deep brain regions [91]. Dondé et al. employed ROAST v3.0 to simulate the electric field induced by tRNS over the left frontotemporal lobe, validating the modulatory effects of high-frequency tRNS in patients with schizophrenia [92]. Similarly, Lewis et al. used SimNIBS to simulate the impact of tDCS on the human blood-brain barrier permeability, ensuring accurate targeting of the current [93]. Beyond using generic models, customized forward models have also been developed to address specific scenarios. For example, Sun et al. investigated the therapeutic potential of tDCS for cortical lesions resulting from decompressive craniectomy (DC). They constructed a finite element head model that included a 12 cm hole at the top of the skull to simulate the surgical imprints [94]. In another study, Breitling et al. created a multimodal head model of a 13-year-old boy using SCIRun, ensuring that the model’s age-specificity aligned with the corresponding subjects [95].
In addition to targeted simulations, forward modeling is extensively used to analyze neuromodulatory mechanisms. For example, Laakso et al. utilized a generic model to compute current pathways and demonstrated that retinal phosphenes could be induced by current [96]. Personalized forward models are increasingly being adopted to simulate more realistic head conditions. This approach involves gathering MRI data from subjects to generate individualized models. Hamajima et al., for instance, used MRI data from 18 healthy subjects to construct forward models for analyzing the effects of tDCS montage-induced polarity changes in lower limb motor regions [97]. In another investigation, Lu et al. developed forward models for healthy older adults and those with mild cognitive impairment (MCI) to simulate the impact of scalp-to-cortex distance on tDCS-induced electric fields [98]. Additionally, Uenishi et al. simulated electric fields induced by prefrontal tDCS in individuals with mood disorders and schizophrenia, finding that the electrophysiological response to stimuli was diminished in those with schizophrenia [99].
Table 3: Applications for forward modeling
Reference Computational Modeling Report Findings
Dondé, C., et al. (2024). Asian journal of psychiatry [92] ROAST hf-tRNS could disrupt the test-retest learning effect in the tone-matching task in individuals with schizophrenia.
Yoon MJ, et al. (2024). Science Report [100] Neurophet tES LAB Optimized tDCS afforded a higher electric field in the target of a stroke patient compared to conventional tDCS.
Lewis, A., et al. (2023). Neuromodulation : journal of the International Neuromodulation Society [93] SimNIBS Although the effects of tDCS were not significant, increases in salivary S100B after a fatiguing cycling task may indicate exercise-induced changes in BBB permeability.
Hamajima H, et al. (2023). Clinical Neurophysiology [97] FreeSurfer Proper montage selection allows reaching deeper regions of the lower-limb motor area with uniform polarization.
Sadeghihassanabadi F, et al. (2022). Journal of Neural Engineering [101] SimNIBS The jaw montage is an optimal choice for maximizing cerebellar stimulation while minimizing unwanted side effects.
Nandi et al. (2022). Brain stimulation [102] SimNIBS The electric field strength calculated by simulation model correlates with a decrease in GABA in the motor cortex.
VanHoornweder et al. (2022). Journal of Neural Engineering [103] SimNIBS Analyzed the impact of individualized versus fixed intensity tES on E-field strength variability.
Uenishi S, et al. (2022). Psychiatry Research: Neuroimaging [99] SimNIBS Attenuated electrophysiological response to transcranial direct current stimulation to the prefrontal cortex in patients with schizophrenia.
Sun W, et al. (2021). Journal of International Medical Research [94] MIMICS A guideline for selecting current and electrode montage settings when performing tDCS on patients after DC.
Lu, H., et al. (2021). Journal of neuroengineering and rehabilitation [98] ROAST Ageing has a prominent, but differential effect on the region-specific SCD and cortical features in older adults with cognitive impairments.
Kasten et al. (2019). Nature Communications [104] ROAST Integrating electric field modeling and neuroimaging to explain inter-individual variability of tACS effects.
Nissim et al. (2019). Frontiers in Aging Neuroscience [105] ROAST Calculating the field distribution by computaional model.
Saturnino et al. (2017). NeuroImage [106] SimNIBS Providing a framework for selective in- and anti-phase TACS of two cortical areas by multiple montages.
Bchinger et al. (2017). The Journal of Neuroscience [17] SimNIBS Assessing whether two stimulation paradigms caused comparable effects by simulation model.
Huang et al. (2017). Brain Stimulation [2] SimNIBS Measured electric potentials intracranially in ten epilepsy patients and estimated electric fields across the entire brain by leveraging calibrated current-flow models.
Breitling, C., et al. (2016). Frontiers in cellular neuroscience [95] SCIRun5 These results suggest that anodal tDCS of the right inferior frontal gyrus could improve interference control in patients with ADHD.
Alekseichuk et al. (2016). Current Biology [107] SimNIBS The anatomical target was validated by simulating the applied electric field.
Laakso, I., et al. (2013). Journal of neural [96] The Duke Model [108] The interference from retinal phosphenes needs to be considered in the design of tACS experiments.
@@ 3 Inverse Optimization for personalized tES
Once the volume conduction model is successfully constructed, a significant concern arises regarding individual variations in tES. In tDCS, due to anatomical differences among individuals, the resulting brain electric field can vary when the same current was administered  [109]. While personalized dosing control can mitigate variations in electric field strength reaching the intended target [110], anatomical variability also presents challenges in determining optimal electrode configurations [35]. Furthermore, tTIS has proven its capacity for achieving focal stimulation [22, 69]. However, the stimulation strategy becomes increasingly complex due to the sophisticated mechanism inherent in tTIS applications [20, 21, 22, 69, 111]. The optimal positioning, size, and spacing of electrodes can significantly amplify the electric field strength within the targeted stimulation zone [112] (Fig. 1 F-I).
@@ 3.1 Formulation of tES optimization problem
Forward modeling enables the estimation of electric field distributions under a given electrode configuration.Based on the simulation results, the effectiveness of stimulation at specific targets can be evaluated (Fig. 1 F). We can use algorithms to identify the optimal configuration for specific targets (Fig. 1 H). The input of the algorithms are objective functions and leadfield matrix. The objective functions incorporate information regarding the target and constraints. The leadfield matrix, or gain matrix, is a matrix that encapsulates the relationship between the stimulation injected at the scalp and the resulting electrical fields in the brain [60]. This matrix is obtained by sequentially applying a unit current through each candidate electrode. Each element of the matrix represents the electrical field in the corresponding area under a given stimulation. By leadfield matrix, we don’t need to run the forward model for each potential configuration, significantly reducing the time cost. The electrical field can be calculated by Eq. (7)
@@
where $s$ represents the index of the activation electrode and the corresponding current intensity, and $L$ denotes the leadfield matrix. The leadfield matrix typically encompasses three dimensions: the electrical channel, mesh volume, and orthogonal direction. The goal of optimization is to maximize or minimize the objective functions, subjecting to constraints, as shown in Eq. (10).
@@
Common goals include maximizing the electric field in the target area [111, 113, 114] and minimizing deviations between the electric field distribution and preset values [114, 115, 116]. Some complex objective functions are discussed in Section 4. And the constraints are primarily designed to ensure the safety of human participants (Fig. 1 G). The total injection current and the maximum individual injection current are denoted as $I_{tot}$ and $I_{ind}$, respectively. For $N$ candidate electrodes, considering the presence of a reference electrode and according to Kirchhoff’s current law, Eq. (11) stipulates that the sum of the absolute values of the current intensities of the candidates and the reference electrode should be less than $2*I_{tot}$. The constraints, Eq. (12) and Eq. (13), are designed to prevent discomfort at the individual electrode interface. The values of $I_{ind}$ and $I_{tot}$ are typically set to 1 mA and 2 mA, respectively. However, these values can vary depending on the specific experimental design. For example, in some protocols, such as ISP stimulation, the amplitude can reach as high as 7 mA [19].
@@
Additionally, constraints can be designed to meet various requirements. Ruffini et al. control the number of active electrodes by penalizing unsuitable solutions [43]. Furthermore, Huang et al. limit the power of the electric field outside the target region to manage the trade-offs between maximal intensity and focality [111].

### v2

@@ 2.3 Simulating Electric Field distribution induced by tES
where $x$, $y$ and $z$ represent three orthogonal directions, and $w_{x}$, $w_{y}$, and $w_{z}$ denote the weights assigned to the respective orthogonal directions. And because tACS usually has low frequency, a quasi-static approximation can be employed to simulate it. In most cases, we can derive the electric field at peak currents from computational models. The temporal variations of the current merely scale the field without altering its distribution within the brain. When multiple channels are out of phase, simulations should be based on temporal dynamics. Saturnino et al. modeled several in-phase and anti-phase tACS scenarios, finding that they produced complex differences in temporospatial stimulation patterns [saturnino2017target].
tTIS employs a distinct mechanism to stimulate the brain. Two-pair tTIS utilizes two pairs of electrodes, while HD-tTIS employs an array of electrodes for stimulation [huang2020optimization]. Both types of stimulation are designed to stimulate the brain by the low frequency envelope formed by high frequency current. The effect of tTIS can be evaluated by the modulation depth, which is defined as the difference between the maximal and minimal value of the envelope. The envelope field with the desired direction can be calculated by the formula proposed by Grossman et al. [grossman2017noninvasive], shown in Eq. (4).
@@
where $\vec{E}_{1}(\vec{r})$ and $\vec{E}_{2}(\vec{r})$ are the first and second distinct fields at the location $\vec{r}$ and $\vec{n}$ is an unit vector along the desired direction. And Huang et al. provided a comprehensive analysis of the equations governing various forms of tTIS [huang2020optimization, huang2019can]. They showed the formulae for the modulation depth along the maximum direction. It is straightforward to find that Eq. (4) is maximized if and only if $\vec{n}$ falls onto the plane spanned by $\vec{E}_{1}$ and $\vec{E}_{2}$. Under such premise, the maximal modulation depth equation can be formulated as follows.
@@ 2.4 Software for Constructing Personalized Head Models
Beyond these general FEM platforms, numerous specialized toolboxes have been developed to streamline various stages of brain stimulation modeling (Table 1). Some, such as FreeSurfer [fischl2012freesurfer], Brainstorm [Tadel2011BrainstormAU], DUNEuro [Medani2023BrainstormDUNEuroAI], Iso2mesh [Tran2020ImprovingMF], SCIRun [dannhauer2012pipeline], COMETS [jung2013comets], and Neurophet tES LAB [karabanov2019can], primarily focus on forward modeling tasks, including anatomical segmentation, mesh generation, and electric field computation. Others, including SimNIBS [saturnino2019simnibs] and ROAST [huang2019realistic], provide end-to-end frameworks that integrate both modeling and optimization within user-friendly environments. Collectively, these tools form a comprehensive computational ecosystem supporting individualized tES research. In the following sections, we introduce the computational frameworks underpinning the major tools listed above, and provide a comparative analysis highlighting their respective strengths, limitations, and suitable application scenarios.
SimNIBS: SimNIBS is an open-source software package for head modeling in tES, supporting both forward simulations and the optimization of stimulation parameters. It enables the construction of individualized head models from MRI data and performs electric field simulations using the FEM. Its strengths include user-friendly interfaces, robust segmentation pipelines, and flexibility in handling multi-channel electrode configurations; however, its performance depends on high-quality MRI data and sufficient computational resources, particularly for complex models.
ROAST: ROAST is an open-source toolbox for personalized head modeling and electric field simulation in MATLAB environments. It leverages SPM12 for tissue segmentation and Iso2mesh for tetrahedral mesh generation, providing anatomically informed modeling of GM, WM, and CSF. ROAST offers a streamlined workflow and supports the processing of MRI data; however, its segmentation accuracy for thin tissue layers can be limited in certain regions.
FreeSurfer: FreeSurfer is a widely used neuroimaging software for cortical surface reconstruction and volumetric segmentation. It provides detailed anatomical labeling of GM and WM, facilitating high-resolution anatomical models suitable for forward modeling of brain activity or electric fields. While FreeSurfer excels in anatomical accuracy, it does not natively perform tES simulations or optimization, although its outputs can be integrated into such pipelines.
Brainstorm: Brainstorm is an open-source platform primarily designed for EEG and MEG source modeling and analysis. It supports head model construction and can interface with FEM solvers for electric field simulations. Its strengths lie in interactive visualization and data‑analysis workflows, while built‑in support for tES optimisation is limited and high‑resolution FEM modeling can demand considerable computation time.
DUNEuro: DUNEuro is an open‑source finite‑element toolbox designed primarily for EEG and MEG forward modelling, with emerging support for tES applications. It provides flexible numerical solvers for complex head geometries and supports integration with custom head models. Its main advantage is high numerical accuracy, though effective use requires familiarity with FEM principles and programming.
Iso2mesh: Iso2mesh is a MATLAB/Octave toolbox for generating tetrahedral and surface meshes from volumetric or surface data. It is compatible with MRI- or CT-derived segmentations, facilitating creation of head‑model meshes that can be used in tES simulation workflows. While flexible and open-source, it is primarily a mesh-generation toolbox and does not itself perform tES simulation or optimization.
SCIRun: SCIRun is a modular problem-solving environment for biomedical simulation and visualization, supporting forward modeling of tES through flexible mesh generation, numerical solvers, and visualization modules. It provides consistent, reusable data structures, simplifying interactive visualization and simulation, but its interface is complex.
COMETS: COMETS is a MATLAB-based platform for modeling tES modalities using individualized or template head models. It allows simulation of electric field distributions for multi-electrode montages and supports protocol design. Its limitations include no built-in MRI segmentation, Windows-only operation, and lack of anisotropic tissue conductivity modeling.
Neurophet tES LAB: Neurophet tES LAB is a commercial platform for modeling and optimizing tES. It integrates head modeling, electric field simulation, and optimization workflows, enabling personalized stimulation protocols. Its main advantage is end-to-end usability, though access is restricted to licensed users.
Table 1: Software and methods for forward modeling and inverse optimization.
Method Scope Website
@@
SimNIBS [saturnino2019simnibs] Both https://simnibs.github.io/simnibs/build/html/index.html
ROAST [huang2019realistic] Both https://github.com/andypotatohy/roast
Brainstorm [Tadel2011BrainstormAU] Modeling https://neuroimage.usc.edu/brainstorm/
@@
DUNEuro [Medani2023BrainstormDUNEuroAI] Modeling https://www.medizin.uni-muenster.de/duneuro/forschung.html
COMSOL [multiphysics1998introduction] Modeling https://www.comsol.com
Genetic Algorithm [stoupis2022non] Optimization https://zenodo.org/records/5907212
@@
MOVEA [wang2022multi] Optimization https://github.com/ncclab-sustech/MOVEA
@@ 3 Inverse Optimization for personalized tES
Once the volume-conduction model has been established, a central issue is inter-individual variability in tES delivery. In tDCS, anatomical differences can cause substantial variation in the induced brain electric field even when identical currents are applied [laakso2015inter]. Personalized dosing helps reduce variability in target-region field strength [evans2020dose], but anatomical heterogeneity also complicates the search for optimal electrode configurations [kasten2019integrating].
For tTIS, relative focality may be achievable, although its extent in humans remains debated. Evidence suggests that apparent focality occurs chiefly when the region of maximal amplitude modulation of the two interfering currents is also the only region effectively modulated [violante2023non, rodgers2025increasing, wessel2023noninvasive]. Moreover, tTIS introduces additional design complexity due to its non-convex properties [grossman2017noninvasive, violante2022non, wang2022multi, huang2019can, huang2018optimized]. In both tDCS and tTIS, electrode geometry—including placement, size, and spacing—can substantially amplify target-region field strength and specificity [caulfield2021optimizing] (Fig. 1F–I).
This section provides an overview of inverse optimization strategies for personalized tES. Section 3.1 introduces the mathematical formulation of the optimization problem. Section 3.2 presents representative algorithmic approaches for optimizing stimulation parameters and configurations.
@@ 3.1 Formulation of tES Optimization Problem
Forward modeling enables the estimation of electric field distributions under a given electrode configuration. Based on the simulation results, the effectiveness of stimulation at specific targets can be evaluated (Fig. 1 F). We can use algorithms to identify the optimal configuration for specific targets (Fig. 1 H). The inputs of the algorithms are objective functions and leadfield matrix. The objective functions incorporate information regarding the target and constraints. The leadfield matrix, or gain matrix, is a matrix that encapsulates the relationship between the stimulation injected at the scalp and the resulting electric fields in the brain [dmochowski2011optimized]. This matrix is obtained by sequentially applying a unit current through each candidate electrode. Each element of the matrix represents the electric field in the corresponding area under a given stimulation. By leadfield matrix, we don’t need to run the forward model for each potential configuration, significantly reducing the time cost. The electric field can be calculated by Eq. (7):
@@
where $s$ represents the index of the activation electrode and the corresponding current intensity, and $L$ denotes the leadfield matrix. The leadfield matrix typically encompasses three dimensions: the electrical channel, mesh volume, and orthogonal direction. The goal of optimization is to maximize or minimize the objective functions, subjecting to constraints, as shown in Eq. (10):
@@
Common goals include maximizing the electric field in the target area [huang2018optimized, guler2016optimization, Dmochowski2011OptimizedMS] and minimizing deviations between the electric field distribution and preset values [Dmochowski2011OptimizedMS, Dmochowski2016OptimalUO, Salman2016ConcurrencyIE]. Some complex objective functions are discussed in Section 4. And the constraints are primarily designed to ensure the safety of human participants (Fig. 1 G). The total injection current and the maximum individual injection current are denoted as $I_{tot}$ and $I_{ind}$, respectively. For $N$ candidate electrodes, considering the presence of a reference electrode and according to Kirchhoff’s current law, Eq. (11) stipulates that the sum of the absolute values of the current intensities of the candidates and the reference electrode should be less than $2*I_{tot}$. The constraints, Eq. (12) and Eq. (13), are designed to prevent discomfort at the individual electrode interface. The values of $I_{ind}$ and $I_{tot}$ are typically set to 1 mA and 2 mA, respectively. However, these values can vary depending on the specific experimental design. For example, in some protocols, such as ISP stimulation, the amplitude can reach as high as 7 mA [voroslakos2018direct]:
@@
Additionally, constraints can be designed to meet various requirements. Ruffini et al. control the number of active electrodes by penalizing unsuitable solutions [ruffini2014optimization]. Furthermore, Huang et al. limit the power of the electric field outside the target region to manage the trade-offs between maximal intensity and focality [huang2018optimized].